English

A problem in Mathematics is given to three students whose chances of solving it are 12,13,14 respectively. If the events of their solving the problem are independent then the probability

Advertisements
Advertisements

Question

A problem in Mathematics is given to three students whose chances of solving it are `1/2, 1/3, 1/4` respectively. If the events of their solving the problem are independent then the probability that the problem will be solved, is ______.

Options

  • `1/4`

  • `1/3`

  • `1/2`

  • `3/4`

MCQ
Fill in the Blanks
Advertisements

Solution 1

A problem in Mathematics is given to three students whose chances of solving it are `1/2, 1/3, 1/4` respectively. If the events of their solving the problem are independent then the probability that the problem will be solved, is `underlinebb(3/4)`.

Explanation:

Let A, B, C be the respective events of solving the problem.

Then, P(A) = `1/2`, P(B) = `1/3` and P(C) = `1/4`.

Here, A, B, C are independent events.

Problem is solved if at least one of them solves the problem.

Required probability is

= P(A ∪ B ∪ C)

= `1 - P(overlineA)P(overlineB)P(overlineC)`

= `1 - (1 - 1/2)(1 - 1/3)(1 - 1/4)`

= `1 - 1/4`

= `3/4`.

shaalaa.com

Solution 2

A problem in Mathematics is given to three students whose chances of solving it are `1/2, 1/3, 1/4` respectively. If the events of their solving the problem are independent then the probability that the problem will be solved, is `underlinebb(3/4)`.

Explanation:

The problem will be solved if one or more of them can solve the problem.

The probability is

`P(Aoverline(BC)) + P(overlineABoverlineC) + P(overline(AB)C) + P(ABoverlineC) + P(AoverlineBC) + P(overlineABC) + P(ABC)`

= `1/2. 2/3. 3/4 + 1/2. 1/3. 3/4 + 1/2 . 2/3. 1/4 + 1/2. 1/3. 3/4 + 1/2. 2/3. 1/4 + 1/2. 1/3. 1/4 + 1/2. 1/3. 1/4`

= `3/4`.

shaalaa.com

Solution 3

A problem in Mathematics is given to three students whose chances of solving it are `1/2, 1/3, 1/4` respectively. If the events of their solving the problem are independent then the probability that the problem will be solved, is `underlinebb(3/4)`.

Explanation:

Let us think quantitively.

Let us assume that there are 100 questions given to A.

A solves `1/2 xx 100` = 50 questions then remaining 50 questions is given to B and B solves `50 xx 1/3` = 16.67 questions.

Remaining `50 xx 2/3` questions is given to C and C solves `50 xx 2/3 xx 1/4` = 8.33 questions.

Therefore, number of questions solved is 50 + 16.67 + 8.33 = 75.

So, required probability is `75/100 = 3/4`.

shaalaa.com
  Is there an error in this question or solution?
2023-2024 (March) Board Sample Paper

Video TutorialsVIEW ALL [2]

RELATED QUESTIONS

A card from a pack of 52 playing cards is lost. From the remaining cards of the pack three cards are drawn at random (without replacement) and are found to be all spades. Find the probability of the lost card being a spade.


Given two independent events A and B such that P (A) = 0.3, P (B) = 0.6. Find 

  1. P (A and B)
  2. P(A and not B)
  3. P(A or B)
  4. P(neither A nor B)

The odds against student X solving a business statistics problem are 8: 6 and odds in favour of student Y solving the same problem are 14: 16 What is the probability that neither solves the problem?


The probability that a student X solves a problem in dynamics is `2/5` and the probability that student Y solves the same problem is `1/4`. What is the probability that

  1. the problem is not solved
  2. the problem is solved
  3. the problem is solved exactly by one of them

A bag contains 3 yellow and 5 brown balls. Another bag contains 4 yellow and 6 brown balls. If one ball is drawn from each bag, what is the probability that, both the balls are of the same color?


Solve the following:

If P(A) = `"P"("A"/"B") = 1/5, "P"("B"/"A") = 1/3` the find `"P"("B'"/"A'")`


The probability of simultaneous occurrence of at least one of two events A and B is p. If the probability that exactly one of A, B occurs is q, then prove that P(A′) + P(B′) = 2 – 2p + q.


10% of the bulbs produced in a factory are of red colour and 2% are red and defective. If one bulb is picked up at random, determine the probability of its being defective if it is red.


If A and B are independent events such that 0 < P(A) < 1 and 0 < P(B) < 1, then which of the following is not correct?


If A and B are independent events such that P(A) = p, P(B) = 2p and P(Exactly one of A, B) = `5/9`, then p = ______.


If A and B′ are independent events then P(A′ ∪ B) = 1 – ______.


Let A and B be two independent events. Then P(A ∩ B) = P(A) + P(B)


Refer to Question 1 above. If the die were fair, determine whether or not the events A and B are independent.


A and B are two events such that P(A) = `1/2`, P(B) = `1/3` and P(A ∩ B) = `1/4`. Find: `"P"("A'"/"B'")`


Let E1 and E2 be two independent events such that P(E1) = P1 and P(E2) = P2. Describe in words of the events whose probabilities are: P1 + P2 – 2P1P2 


Two dice are tossed. Find whether the following two events A and B are independent: A = {(x, y): x + y = 11} B = {(x, y): x ≠ 5} where (x, y) denotes a typical sample point.


If two events are independent, then ______.


Let P(A) > 0 and P(B) > 0. Then A and B can be both mutually exclusive and independent.


If A and B are mutually exclusive events, then they will be independent also.


If A and B are two events such that P(A|B) = p, P(A) = p, P(B) = `1/3` and P(A ∪ B) = `5/9`, then p = ______.


If A, B are two events such that `1/8 ≤ P(A ∩ B) ≤ 3/8` then


Given two events A and B such that (A/B) = 0.25 and P(A ∩ B) = 0.12. The value P(A ∩ B') is ______.


What are independent events in probability?


Which is the equivalent and most commonly used test for independent events E and F?


A die is thrown once. What is the sample space \[S\]?


A die is thrown once. If E is the event that the number appearing is a multiple of 3, which set represents E?


For \[E=\{3,6\}\] when a die is thrown once, what is \[P(E)\]?


For \[P(E)=\frac{1}{3}\] and \[P(F)=\frac{1}{2}\], what is \[P(E)P(F)\]?


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×